OpenStudy (anonymous):

a and b are odd integers and c is an even integer, which of the following would be an odd integer? a + b + c ab – bc a x b x c c(a + b)

OpenStudy (praxer):

i suppose b if a=7,b=5 and c=2 then then it becomes 7*5-5*2= 35-10=25

OpenStudy (praxer):

check it again....

OpenStudy (anonymous):

I thought it was C @praxer

OpenStudy (praxer):

but then it comes 7*5*2=70 which is even.. so it is B